# version 1.0

### v6

@@ 5.2 Optimal Solution of the LQR
We know from Chapter 3 that the LQR is a special case of a finite-time Markov decision process and hence admits an optimal policy. However, the optimal policy is in general a mapping from states to actions, both of which are infinite Euclidean spaces in our case.
However, it can be shown that the optimal policy for an LQR is a linear policy, that is, a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$! This astonishing fact, together with a remarkable property of the value function, is formally given in the following theorem. We state the theorem for a finite number of iterations and for a linear time-invariant system. Generalizing this to a time-varying linear system is left as exercise 5.6.
@@
For the LQR problem with finite $T>0$ iterations, there exist positive semi-definite matrices $S_{t}$, ${K}_{t}$ and scalars $c_{t}$, such that the optimal value function and optimal policy satisfy:
$\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\top}{S}_{t}\mathbf{x}+c_{t},\quad\mathbf{u}_{t}^{*}(\mathbf{x})={K}_{t}\mathbf{x}.$
@@
Recall that the Bellman Optimality equation for a general MDP is written as
$\mathbf{v}_{t}^{*}(s)=\max_{a}\left\{R_{sa}+\sum_{s^{\prime}}P^{a}_{ss^{\prime}}\mathbf{v}_{t+1}^{*}(s^{\prime})\right\}.$
We henceforth write the special case applicable to LQR, and prove the theorem by backward induction on $t$.
@@
$\mathbf{v}^{*}_{T}(\mathbf{x})=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}\big\}.$
Since ${R}\succ 0$, a minimum occurs at $\mathbf{u}=0$. Hence, ${S}_{T}={Q}\succ{0}$, ${K}_{T}={0}$ and $c_{T}=0$.
Inductive Step: Assume for some $t>1$ that $\mathbf{u}_{t}^{*}(\mathbf{x})={K}_{t}\mathbf{x}$ and $\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\intercal}{S}_{t}\mathbf{x}+c_{t}$. Then we have
$\mathbf{v}^{*}_{t-1}(\mathbf{x})$ $=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+\mathop{\mbox{\bf E}}_{\mathbf{w}}\mathbf{v}^{*}_{t}({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})\big\}$
$=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+\mathop{\mbox{\bf E}}_{\mathbf{w}}\left[({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})^{\intercal}S_{t}({A}\mathbf{x}+{B}\mathbf{u}+\mathbf{w})\right]\big\}+c_{t}$
$=\min_{\mathbf{u}}\big\{\mathbf{x}^{\intercal}{Q}\mathbf{x}+\mathbf{u}^{\intercal}{R}\mathbf{u}+({A}\mathbf{x}+{B}\mathbf{u})^{\intercal}{S}_{t}({A}\mathbf{x}+{B}\mathbf{u})+\mathop{\mbox{\bf E}}_{\mathbf{w}}[\mathbf{w}^{\top}S_{t}\mathbf{w}]\big\}+c_{t},$
where in the last equality we used the fact that $\mathop{\mbox{\bf E}}[\mathbf{w}]=0$, and that the perturbation is independent of the state and control. We henceforth denote the second moment of the noise by $\mathop{\mbox{\bf E}}[\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}I$. This implies that for any matrix $M$ which is independent of $\mathbf{w}$, we have $\mathop{\mbox{\bf E}}[\mathbf{w}^{\top}M\mathbf{w}]=\mathop{\mbox{\bf E}}[M\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}\cdot M\bullet I=\sigma^{2}{\bf Tr}(M)$.
Since $R$ and $S_{t}$ are positive semidefinite, the above expression is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
$\nabla_{\mathbf{u}}\Big[{\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{u}^{\intercal}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\intercal}S_{t}(A\mathbf{x}+B\mathbf{u})+\sigma^{2}{\bf Tr}(S_{t})}]\Big]=0.$
Since $Q,R,S_{t}$} are all positive semi-definite and symmetric, this gives us
$2R\mathbf{u}+2B^{\intercal}S_{t}A\mathbf{x}+2B^{\intercal}S_{t}B\mathbf{u}=0\implies(R+B^{\intercal}S_{t}B)\mathbf{u}=-B^{\intercal}S_{t}A\mathbf{x}$}.
Hence, we have that
$\mathbf{u}_{t-1}(\mathbf{x})=K_{t-1}\mathbf{x},\ \ \ K_{t-1}=-(R+B^{\intercal}S_{t}B)^{-1}(B^{\intercal}S_{t}A)$}.
Notice that the matrix $R+B^{\top}S_{t}B$ is full rank since $R\succ 0$. Plugging this in the expression for the value function, we get
$\mathbf{v}_{t-1}^{*}(\mathbf{x})$
$={\mathbf{x}^{\intercal}Q\mathbf{x}+\mathbf{x}^{\intercal}K_{t-1}^{\intercal}RK_{t-1}\mathbf{x}+\mathbf{x}^{\intercal}(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})\mathbf{x}+\sigma^{2}{\bf Tr}(S_{t})}+c_{t}$
$=\mathbf{x}^{\intercal}{S}_{t-1}\mathbf{x}+c_{t-1}$
where $S_{t-1}=Q+K_{t-1}^{\intercal}RK_{t-1}+(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})$} is positive definite, and $c_{t-1}=c_{t}+\sigma^{2}{\bf Tr}(S_{t})$.
@@ 5.3 Infinite Horizon LQR
Although infinite horizon problems can never appear in the real world, in practice the horizon length may be unknown. Although there are techniques to iteratively update the horizon length with geometrically increasing guesstimates, these are slow and impractical.
An alternative is to solve the LQR for the infinite horizon case, which results in a very elegant solution. Not only that, this turns out to be a practical solution to which the finite-horizon solution converges quickly.
The infinite horizon LQR problem can be written as follows:
$\min_{\mathbf{u}(\mathbf{x})}\lim_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\right]$
$\text{ s.t. }\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t},$
where similar to the finite horizon case, we assume that the perturbations $\mathbf{w}_{t}$ are i.i.d. zero mean and have bounded variance. We henceforth assume that the perturbations are all zero for simplicity, and adding zero-mean bounded-variance perturbations is left as exercise 5.6.
Without perturbations, the Bellman optimality equation becomes
@@
Assuming that the value function is quadratic, we have:
@@
Similarly to the finite-time case, we have the following result.
Lemma 5.3.
The optimal policy is given by a transformation $\mathbf{u}^{*}(\mathbf{x})={K}\mathbf{x}$.
Proof.
At optimality, the gradient with respect to $\mathbf{u}$ of the above equation is zero. This implies that
$2R\mathbf{u}+2B^{\top}SB\mathbf{u}+2B^{\top}SA\mathbf{x}=0$},
which implies
$K=-(R+B^{\top}SB)^{-1}(B^{\top}SA)$}.
Plugging this back into the Bellman equation, we obtain
@@
Since this equality holds for all $\mathbf{x},$, we get
@@
Simplifying this expression we obtain
$S=Q+A^{T}SA-A^{T}SB\left({R}+B^{T}SB\right)^{-1}B^{T}SA.$
This equation is called the Discrete Time Algebraic Riccati Equation (DARE) and is one of the most important equations in control theory.
@@ 5.3.1 Solving the DARE
In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by $R+B^{\top}SB$} to form a quadratic equation in S{S}. Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial in the one-dimensional case. The reason we only consider the nonnegative root is that the value function is nonnegative by our choice of quadratic convex costs.
In higher dimension, the correct generalization is somewhat more difficult to see, but using Lemma 4.5, the equation can be reformulated as the solution to the following semidefinite program.
$\min_{{X}\in\mathcal{S}^{d_{x}\times d_{x}}}\langle{X,I}\rangle$
$\text{subject to}\begin{bmatrix}{X}&B^{T}SA\\ A^{T}SB&{R}+B^{T}SB\end{bmatrix}\succeq 0,$
$\text{and }S=Q+A^{T}SA-{X}.$
A popular alternative is to adapt the value iteration algorithm to this setting by iteratively applying the right side of the aforementioned equation as an operator. Concretely, starting from ${S}_{0}\succeq 0$, we choose
$S_{t+1}=Q+A^{T}S_{t}A-A^{T}S_{t}B\left({R}+B^{T}S_{t}B\right)^{-1}B^{T}S_{t}A.$
@@ 5.4 Robust Control
One of the most important and prominent generalizations is that of robust control, also known as $\text{H}_{\infty}$-control. The goal of $\text{H}_{\infty}$-control is to come up with a more robust solution, by considering the optimal solution vs. any possible perturbation sequence from a certain family. Formally, our aim is to solve the following:
$\begin{array}[]{c}\min_{\mathbf{u}_{1:T}(\mathbf{x})}\max_{\mathbf{w}_{1:T}}\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}{Q}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}{R}\mathbf{u}_{t}\\ \text{ s.t. }\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},\end{array}$
where we use the exact same notation as in previous sections.
Notice that this problem does not fall into the category of a Markov Decision Process, since $\mathbf{w}_{t}$’s are no longer stochastic. Therefore, the typical Bellman optimality criterion does not apply.
Instead, the appropriate techniques to address this problem come from the theory of dynamical games. The above is, in fact, a two-player zero sum game. The first player tries to minimize the costs, and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
Using techniques from dynamic games, the following can be shown. References are provided in the bibliographic section.
@@
For the $\text{H}_{\infty}$ control problem with quadratic convex loss functions, the optimal solution is a linear policy of the state given by
$\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}.$
@@ 5.5 Bibliographic Remarks
While seemingly a natural corollary of equivalent results for MDPs, the existence of value functions for linear control is made subtle by the continuous nature of the state space and stabilizability, or the lack thereof. A discussion of sufficient and necessary conditions for the existence of solutions to Riccati equations can be found in [Ber07, HSK96].
The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to misspecification of system parameters. This suggests that robustness specifications constitute yet another axis of performance for control algorithms and may at times be at odds with the objective of optimal control. For in-depth treatment of $H_{\infty}$ control, see [ZDG96b]. The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary; see also [Tu18] for a recent exposition.

### v8

@@ 5.2 Optimal Solution of the LQR
We know from Chapter 3 that the LQR is a special case of a finite-horizon Markov decision process and hence admits an optimal policy. However, the optimal policy is in general a mapping from states to actions, both of which are infinite Euclidean spaces in our case.
It turns out that the optimal policy for an LQR is a linear policy, that is, a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$. This remarkable fact, together with a corresponding quadratic form for the value function, is given in the following theorem. We state the theorem for a finite horizon and a linear time-invariant system. Generalizing this to a time-varying linear system is left as Exercise 5.6.
@@
For the finite-horizon LQR problem, there exist symmetric matrices $S_{t}\succeq 0$, gains $K_{t}\in{\mathbb{R}}^{d_{u}\times d_{x}}$, and scalars $c_{t}\in{\mathbb{R}}$ such that, for every stage $t$,
$\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\top}S_{t}\mathbf{x}+c_{t},\qquad\mathbf{u}_{t}^{*}(\mathbf{x})=K_{t}\mathbf{x}.$
@@
For $t=1,\dots,T$, let $\mathbf{v}_{t}^{*}(\mathbf{x})$ denote the optimal cost-to-go from stage $t$ onward when $\mathbf{x}_{t}=\mathbf{x}$, and set $\mathbf{v}_{T+1}^{*}(\mathbf{x})\equiv 0$. We prove the claim by backward induction on $t$.
@@
$\mathbf{v}_{T}^{*}(\mathbf{x})=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}\right\}.$
Since $R\succ 0$, the unique minimum occurs at $\mathbf{u}=0$. Hence
$S_{T}=Q,\qquad K_{T}=0,\qquad c_{T}=0.$
Inductive Step: Assume for some $t>1$ that
$\mathbf{v}_{t}^{*}(\mathbf{x})=\mathbf{x}^{\top}S_{t}\mathbf{x}+c_{t}$
for a symmetric matrix $S_{t}\succeq 0$. Then
$\mathbf{v}_{t-1}^{*}(\mathbf{x})$ $=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+\mathbb{E}_{\mathbf{w}}\!\left[\mathbf{v}_{t}^{*}(A\mathbf{x}+B\mathbf{u}+\mathbf{w})\right]\right\}$
$=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+\mathbb{E}_{\mathbf{w}}\!\left[(A\mathbf{x}+B\mathbf{u}+\mathbf{w})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u}+\mathbf{w})\right]\right\}+c_{t}.$
Using $\mathbb{E}[\mathbf{w}]=0$ and $\mathbb{E}[\mathbf{w}\mathbf{w}^{\top}]=\Sigma$, we obtain
$\mathbb{E}_{\mathbf{w}}\!\left[(A\mathbf{x}+B\mathbf{u}+\mathbf{w})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u}+\mathbf{w})\right]=(A\mathbf{x}+B\mathbf{u})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u})+{\bf Tr}(S_{t}\Sigma).$
Therefore,
$\mathbf{v}_{t-1}^{*}(\mathbf{x})=\min_{\mathbf{u}}\left\{\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{u}^{\top}R\mathbf{u}+(A\mathbf{x}+B\mathbf{u})^{\top}S_{t}(A\mathbf{x}+B\mathbf{u})\right\}+c_{t}+{\bf Tr}(S_{t}\Sigma).$
Since $R\succ 0$ and $S_{t}\succeq 0$, the expression inside the braces is a strictly convex quadratic function of $\mathbf{u}$. Its gradient is
$2(R+B^{\top}S_{t}B)\mathbf{u}+2B^{\top}S_{t}A\mathbf{x}.$
Setting the gradient to zero yields
$(R+B^{\top}S_{t}B)\mathbf{u}=-B^{\top}S_{t}A\mathbf{x}.$
Since $R\succ 0$, the matrix $R+B^{\top}S_{t}B$ is positive definite and therefore invertible. Hence the optimal control is
$\mathbf{u}_{t-1}^{*}(\mathbf{x})=K_{t-1}\mathbf{x},\qquad K_{t-1}=-(R+B^{\top}S_{t}B)^{-1}B^{\top}S_{t}A.$
Substituting this back into the value function gives
$\mathbf{v}_{t-1}^{*}(\mathbf{x})$ $=\mathbf{x}^{\top}Q\mathbf{x}+\mathbf{x}^{\top}K_{t-1}^{\top}RK_{t-1}\mathbf{x}$
$\qquad\qquad+\mathbf{x}^{\top}(A+BK_{t-1})^{\top}S_{t}(A+BK_{t-1})\mathbf{x}+c_{t}+{\bf Tr}(S_{t}\Sigma)$
$=\mathbf{x}^{\top}S_{t-1}\mathbf{x}+c_{t-1},$
where
$S_{t-1}=Q+K_{t-1}^{\top}RK_{t-1}+(A+BK_{t-1})^{\top}S_{t}(A+BK_{t-1}),$
and
$c_{t-1}=c_{t}+{\bf Tr}(S_{t}\Sigma).$
Since each term in the definition of $S_{t-1}$ is positive semidefinite, we have $S_{t-1}\succeq 0$. This completes the induction. ∎
@@ 5.3 Infinite Horizon LQR
Although real control tasks always have finite duration, in practice the horizon length may be very large or unknown. This motivates the infinite-horizon formulation, which leads to an especially elegant stationary controller.
In this section we consider the deterministic infinite-horizon total-cost problem
$\min_{\pi}\quad$ $\sum_{t=1}^{\infty}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}R\mathbf{u}_{t}$
s.t. $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t},$
where $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$. We assume throughout that $Q\succ 0$, $R\succ 0$, and that the pair $(A,B)$ is stabilizable.
Persistent additive noise is more naturally treated with an average-cost formulation rather than a total-cost one; see Exercise 5.6.
Under the above assumptions, the optimal policy is stationary and linear, and the value function has quadratic form. Writing the Bellman optimality equation for the total cost, we obtain
@@
Assuming that the value function is quadratic, say $\mathbf{v}^{*}(\mathbf{x})=\mathbf{x}^{\top}S\mathbf{x}$, we get
@@
Since the objective is strictly convex in $\mathbf{u}$, the optimum is characterized by a vanishing gradient:
$2R\mathbf{u}+2B^{\top}SB\mathbf{u}+2B^{\top}SA\mathbf{x}=0.$
Hence the optimal policy is linear and given by
$\mathbf{u}^{*}(\mathbf{x})=K\mathbf{x},\qquad K=-(R+B^{\top}SB)^{-1}B^{\top}SA.$
Plugging this back into the Bellman equation yields
@@
Since this equality holds for all $\mathbf{x}$, we obtain
@@
Simplifying this expression gives
$S=Q+A^{\top}SA-A^{\top}SB\left(R+B^{\top}SB\right)^{-1}B^{\top}SA.$
This equation is called the Discrete-Time Algebraic Riccati Equation (DARE) and is one of the most important equations in control theory. Its stabilizing solution determines the optimal infinite-horizon controller.
@@ 5.3.1 Solving the DARE
In the one-dimensional case, although the right-hand side of the equation is a rational function, we can multiply both sides by $R+B^{\top}SB$ to form a quadratic equation in $S$. Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial. We only consider the nonnegative root because the value function is nonnegative by our choice of quadratic convex costs.
In higher dimensions, a convenient convex formulation is obtained from the associated Riccati inequality. By the Schur complement, the matrix inequality
$\begin{bmatrix}Q+A^{\top}SA-S&A^{\top}SB\\ B^{\top}SA&R+B^{\top}SB\end{bmatrix}\succeq 0$
is equivalent to
$S\preceq Q+A^{\top}SA-A^{\top}SB\left(R+B^{\top}SB\right)^{-1}B^{\top}SA.$
Under the standard stabilizability assumptions, the maximal positive semidefinite solution of this Riccati inequality is precisely the stabilizing solution of the DARE. Therefore, one can recover $S$ via the semidefinite program
$\max_{S\in\mathcal{S}^{d_{x}\times d_{x}}}\langle S,I\rangle$
subject to
$\begin{bmatrix}Q+A^{\top}SA-S&A^{\top}SB\\ B^{\top}SA&R+B^{\top}SB\end{bmatrix}\succeq 0,\qquad S\succeq 0.$
A popular alternative is the Riccati iteration: starting from some $S_{0}\succeq 0$, define
$S_{t+1}=Q+A^{\top}S_{t}A-A^{\top}S_{t}B\left(R+B^{\top}S_{t}B\right)^{-1}B^{\top}S_{t}A.$
Under the same assumptions, this sequence converges to the stabilizing solution of the DARE.
@@ 5.4 Robust Control
One of the most important and prominent generalizations is robust control, also known as $H_{\infty}$-control. The goal of $H_{\infty}$-control is to come up with a more robust solution by comparing the controller against an adversary that chooses the disturbance sequence.
A standard finite-horizon formulation fixes a robustness parameter $\gamma>0$ and considers
$\begin{array}[]{cl}\min_{\pi_{1},\dots,\pi_{T}}\max_{\psi_{1},\dots,\psi_{T}}&\displaystyle\sum_{t=1}^{T}\left(\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}R\mathbf{u}_{t}-\gamma^{2}\mathbf{w}_{t}^{\top}\mathbf{w}_{t}\right)\\[10.00002pt] \text{s.t.}&\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},\\ &\mathbf{u}_{t}=\pi_{t}(\mathbf{x}_{t}),\qquad\mathbf{w}_{t}=\psi_{t}(\mathbf{x}_{t}).\end{array}$
The parameter $\gamma$ controls the tradeoff between nominal performance and robustness.
Notice that this problem does not fall into the category of a Markov decision process, since the disturbances are no longer stochastic. Instead, the appropriate techniques come from the theory of dynamic games. The above problem is a two-player zero-sum game: the controller tries to minimize the cost, while the adversary tries to maximize it, both subject to the same dynamical constraint.
Using techniques from dynamic games, one can show the following. References are provided in the bibliographic section.
@@
Under the standard solvability conditions for finite-horizon $H_{\infty}$ control, the above dynamic game admits a saddle point in linear state-feedback strategies. In particular, there exist matrices $K_{t}$ and $L_{t}$ such that
$\mathbf{u}_{t}=K_{t}\mathbf{x}_{t},\qquad\mathbf{w}_{t}=L_{t}\mathbf{x}_{t}.$
@@ 5.5 Bibliographic Remarks
While the finite-horizon derivation above follows the spirit of dynamic programming for Markov decision processes, the infinite-horizon theory is subtler because of the continuous nature of the state space and the role of stabilizability. A discussion of sufficient and necessary conditions for the existence of stabilizing solutions to Riccati equations can be found in [BER07, HSK96].
The development of $H_{\infty}$ control was driven in part by the observation, made in the seminal paper of Doyle [DOY78], that optimal controllers need not be robust to misspecification of system parameters. This suggests that robustness specifications constitute yet another axis of performance for control algorithms and may at times be at odds with the objective of optimal control. For in-depth treatment of $H_{\infty}$ control, see [ZDG96b]. The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary; see also [TU18] for a recent exposition.